Qd = 90 – 6P


Qs = - 20 + 4P


Calculate consumer surplus and producer surplus at the equilibrium price?

1
Expert's answer
2022-02-21T14:42:11-0500

"Q_d= 90- 6P"

"Q_s= -20+4P"

At Equilibrium, "Q_d= Q_s"

"90-6p= -20+4P"

"10P= 110"

P= 11

"Q^*=90- 6(11)= 24"

Y intercept P"_{max}= Q=0" for demand curve

0=90-6P

P"_{max}" = 15

Y intercept for supply curve, P"_{min}" = When Q= 0

0= -20+4P

P= 5


"Consumer surplus= \frac {1}{2}\times Q_d \times \Delta P"

"\frac{1}{2}\times 24 \times (15-11)"

= 48


Producer Surplus= "\frac{1}{2}\times 24 \times (11-5)"

"= 72"